9/5a - 1/2 = -6 3/8
Found 2 solutions by lynnlo, stanbon:
Answer by lynnlo(4176)
Answer by stanbon(75887)

9/5a - 1/2 = -6 3/8
----
(9/5)a - (1/2) = (-51/8)
----
Multiply thru by 40 to get:
72a - 20 = -255
72a = -235
-----------
a = -3.2639
=======================
Cheers,
Stan H.
============================